OpenStudy (anonymous):

(x^5+18x^2-27x)/(x+3)

OpenStudy (mimi_x3):

\[\frac{x^{5}+18x^{2}-27x}{x+3} \] what do you want to do with it? simplify?

OpenStudy (anonymous):

find the quotient

sam (.sam.):

Synthetic division

OpenStudy (anonymous):

idk it says find the quotient and enter your answer n the decending order below.

OpenStudy (mimi_x3):

would long division work? i havent done polynomials in a while..

OpenStudy (anonymous):

i think so

OpenStudy (mimi_x3):

I think that this is easier; not sure if it would work: \(P(x)=D(x)*Q(x)\)

sam (.sam.):

You do it like the pic below, \[x^4-3x^3+9x^2-9x\] is your quotient

OpenStudy (shubhamsrg):

num is(x^5+18x^2-27x) x^5 -3x^4 + 3x^4 + 9x^3 - 9x^3 +27x^2 -9x^2 -27x => x^5 - 3x^4 + 9x^3 -9x^2 + 3x^4 -9x^3 + 27x^2-27x =>x(x^4 -3x^3 + 9x^2-9x) + 3(x^4 -3x^3 + 9x^2-9x) =>(x^4 -3x^3 + 9x^2-9x)(x+3) which gives your ans and by thew way,,long division will also work..

sam (.sam.):

@Mimi_x3 there are other ways, since the remainder is 0, we can say that x=-3 is a factor of (x^5+18x^2-27x) then f(-3)=x^5+18x^2-27x

sam (.sam.):

but that's finding zero , lol

sam (.sam.):

just use the synthetic division ;)

OpenStudy (anonymous):

is it easier

OpenStudy (mimi_x3):

its the same thing i suppose..

OpenStudy (anonymous):

okaii